### v4

# Life Finds a Way: Emergence of Cooperative Structures in Adaptive Threshold Networks

Abstract
There has been a long debate on how new levels of organization have evolved [20]. The evolution of higher-level organization can appear unlikely, since cooperation must prevail over competition [22]. One well-studied example is the emergence of autocatalytic sets [9], which are often considered a prerequisite for the evolution of life.
Using a simple model, we investigate how varying bias toward cooperation versus antagonism shapes network dynamics, revealing that higher-order organization emerges even amid pervasive antagonistic interactions. In general, we observe that a quantitative increase in the number of elements in a system leads to a qualitative transition.
We present a random threshold-directed network model [24] that integrates node-specific traits with dynamic edge formation and node removal, simulating arbitrary levels of cooperation and competition. In our framework, intrinsic node values determine directed links through various threshold rules. Our model generates a multi-digraph with signed edges (reflecting support/antagonism, labeled “help”/“harm”), which ultimately yields two parallel yet interdependent threshold graphs. Incorporating temporal growth and node turnover in our approach allows exploration of the evolution, adaptation, and potential collapse of communities and reveals regime changes in both connectivity and resilience.
Our findings extend classical random threshold and Erdős-Rényi models [8], offering new insights into adaptive systems in biological and economic contexts, with emphasis on the application to Collective Affordance Sets [18]. This framework will also be useful for making predictions that will be tested by ongoing experiments of microbial communities in soil.
keywords:
adaptive threshold networks , cooperation under antagonism , strongly connected components , regime change , crossover , collective affordance sets , emergent organization , complex systems biology , microbial communities
journal: Theoretical Biology
[…]
[label1]organization=SUNY Finger Lakes CC, addressline=3325 Marvin Sands Dr, city=Canandaigua, postcode=14424, state=NY, country=USA
[…]
[label2]organization=SUNY Binghamton, addressline=4400 Vestal Pkwy E, city=Binghamton, postcode=13902, state=NY, country=USA
[…]
[label3]organization=Institute for Systems Biology, addressline=401 Terry Ave N, city=Seattle, postcode=98109, state=WA, country=USA
{graphicalabstract}[Uncaptioned image]
{highlights}
Even in antagonistic environments, the system finds “order for free”: a mutually enabling SCC self-assembles and persists.
When diversity and binding chance cross a threshold, quantity gives way to quality: a giant SCC takes shape.
1 Introduction
Life drives — and is driven by — interactions. Each organism brings a unique and vast combination of causal properties into its environment, continuously adapting and “tinkering” to establish beneficial interactions or mitigate harmful ones. Despite evolutionary pressures and competition, we regularly observe that complex, self-sustaining networks of mutual support routinely emerge in living systems [22]. Understanding how and why such networks reliably form is central to both theoretical biology and complex systems science.
Microbial communities illustrate this phenomenon particularly well. Despite (or, possibly due to) immense biochemical and ecological diversity, stable interdependent communities arise repeatedly. Even when initial interactions are predominantly antagonistic, microbes find ways to coexist or even benefit each other, forming what can be viewed as ”collective affordance sets” [18]: jury-rigged, self-organized structures that exploit combinations of causal properties among community members.
To empirically examine this phenomenon, ongoing experimental work aims to mix 56 bacterial and 56 fungal species, to observe if and how novel stable ecosystems emerge from a huge potential space of interaction patterns. While our primary interest is in complex soil communities of microorganisms, this concept extends beyond biology. Human economies share remarkably similar dynamics. Technological innovation and market diversity are driven by recombinations of existing products, services, and ideas. As more firms or inventors enter the market, new products and services find ”niches,” creating a self-reinforcing and evolving web of interdependence.
Cooperative and interdependent structures in microbial communities need not arise from explicitly altruistic strategies. A prominent alternative framework is the Black Queen Hypothesis, which explains the emergence of metabolic dependencies through adaptive gene loss and the retention of “leaky” functions by a subset of community members. In this view, organisms may reduce individual costs by abandoning costly functions while relying on others to supply shared resources, producing cooperation-like patterns without cooperative intent. The present work is not intended to model gene loss or metabolic specialization directly. Rather, it focuses on the dynamical consequences of structured help and harm interactions at the network level, abstracting away from the evolutionary pathways by which such interactions arise. As such, the framework developed here should be viewed as complementary to Black Queen–type mechanisms, addressing how interdependence (once present) shapes the stability and organization of communities.
In this paper, we present an adaptive threshold network model with simple assumptions, designed to explore the conditions that foster stable, cooperative structures. Rather than focus on specific evolutionary games or particular ecological scenarios, our model investigates the broader space of possible interaction patterns, identifying conditions under which large, strongly connected communities (SCCs) reliably form and persist. Moreover, because “alternative community states can arise as a consequence of system dynamics without being driven by environmental differences” [11], the broad, environment-agnostic design of our model is well justified.
Using a simple model, our results demonstrate that networks of mutual support robustly emerge even when antagonistic interactions predominate, provided certain thresholds of cooperation and species longevity are met. Moreover, these emergent communities exhibit distinctive patterns, such as stable configurations of key cooperative nodes and systematic turnover influenced by antagonism. Thus, our findings offer insights not only into the formation of biological ecosystems but also into analogous processes in economics and social systems, highlighting universal principles underlying collective organization and adaptive complexity.
2 Model
2.0.1 Model definition (birth–binding–culling dynamics)
We simulate an evolving signed directed multigraph $G_{t}=(V_{t},E_{t})$ using NetworkX’s MultiDiGraph, where vertices represent agents and directed edges represent directed interactions. Time is discrete, $t=1,2,\dots$, and each time-step consists of: (i) adding one new node, (ii) probabilistically forming directed help/harm edges between the newcomer and all incumbents, (iii) aging all nodes by one tick, and (iv) removing any nodes that become eligible for deletion.
Parameters.
The dynamics are governed by the following global parameters:
  • 1.
    $\lambda>0$: Poisson rate controlling heterogeneity in opportunities for pairwise interaction.
  • 2.
    $p_{\max}\in(0,1]$: baseline upper bound used to convert opportunity into an effective binding threshold/probability.
  • 3.
    $\rho\in[0,1]$: global help–harm mixture; $\rho$ is the probability that a successful binding event is harmful (and $1-\rho$ that it is helpful).
  • 4.
    $m_{\max}\in\mathbb{N}$: maximum number of parallel edges allowed per ordered pair and sign (we use $m_{\max}=6$ unless stated otherwise).
  • 5.
    $L_{S}\in\mathbb{N}$: minimum age before a node becomes eligible for removal (a lifespan/juvenile protection period, $L_{S}=100$ by default).
  • 6.
    $\eta\in[0,1]$: removal threshold for the fraction of incoming harm (we use $\eta=0.5$ unless stated otherwise).
Node traits (intrinsic capacities).
When a new node $i$ enters, it is endowed with four independent traits drawn i.i.d. uniformly from $[0,1]$:
  • 1.
    $h_{\text{in}}(i)$: receptivity to beneficial (help) interactions,
  • 2.
    $h_{\text{out}}(i)$: capacity to provide beneficial (help) interactions,
  • 3.
    $\gamma_{\text{in}}(i)$: vulnerability to detrimental (harm) interactions,
  • 4.
    $\gamma_{\text{out}}(i)$: capacity to inflict detrimental (harm) interactions.
We treat help and harm as independent dimensions rather than opposite ends of a single axis. Biologically, an organism can plausibly both support mutualists and inhibit competitors via distinct pathways, so the ability to help need not constrain the ability to harm (and vice versa).
Pair-specific binding opportunity and effective threshold.
When node $i$ enters at time $t$, it “tests” every incumbent node $j\in V_{t-1}$ in both directions $(i\to j)$ and $(j\to i)$. For each ordered pair $(u,v)\in\{(i,j),(j,i)\}$ we draw
$k_{uv}\sim\mathrm{Poisson}(\lambda),\qquad P_{uv}=\min\bigl(1,\;k_{uv}\,p_{\max}\bigr).$
Here $k_{uv}$ represents the (random) number of opportunities for $u$ to interact with $v$ during this time-step, and $P_{uv}$ converts that opportunity into an effective binding scale capped at $1$.
Edge formation (help vs harm, with multiplicity).
For each ordered pair $(u,v)$ involving the newcomer and an incumbent, we perform up to $m_{\max}$ binding attempts. Each attempt proceeds as follows:
  1. 1.
    Draw the interaction type: harm with probability $\rho$, help with probability $1-\rho$.
  2. 2.
    Conditional on the type, form a directed edge if the relevant trait mismatch is within the effective threshold:
    $\text{help:}\quad|h_{\text{out}}(u)-h_{\text{in}}(v)|<P_{uv},\qquad\text{harm:}\quad|\gamma_{\text{out}}(u)-\gamma_{\text{in}}(v)|<P_{uv}.$
Successful attempts add a parallel directed edge of the corresponding sign; unsuccessful attempts add nothing. This allows repeated affordances or repeated adverse interactions between the same ordered pair while preventing unbounded multiplicity.
Aging and culling (node removal rule).
After binding is evaluated for the new node, every node ages by one tick. A node $x$ becomes eligible for removal only once its age is at least $L_{S}$. Eligible nodes are removed (along with all incident edges) if either:
  1. 1.
    $x$ has no incoming edges (no support or influence from others), or
  2. 2.
    the incoming harm fraction exceeds $\eta$:
    $\frac{\mathrm{harm}_{\mathrm{in}}(x)}{\mathrm{help}_{\mathrm{in}}(x)+\mathrm{harm}_{\mathrm{in}}(x)}>\eta.$
Edges do not decay on their own; they disappear only when one endpoint is deleted. Repeating this birth–binding–culling cycle yields the evolving signed network used in subsequent experiments. We use $\eta=0.5$ unless specified otherwise.
AABB
AABBCC
AABBCCDD
AABBCCDDEE
AABBDDEE
Figure 1: Sample simulation showing graph generation and node removal. Top row (left to right): Steps 2–4. Bottom row: Step 5 (E added, C marked for removal) and Step 6 (C and its edges removed). For clarity, the lifespan condition is omitted; in the full model, a node is only eligible for removal once its age is at least $L_{S}$.
2.0.2 Sample Simulation
Step 1: Node A is added.
Step 2: Node B added, harm (red) edge added $B\to A$ as $|\gamma_{\text{in},A}-\gamma_{\text{out},B}|<P_{BA}$, and help (blue) edge added $A\to B$ as $|h_{\text{out},A}-h_{\text{in},B}|<P_{AB}$.
Step 3: Node C added, 3 blue and 2 red links added after threshold test between C and existing nodes A,B.
Step 4: Node D added, 2 blue links added after threshold test between D and existing nodes A,B,C.
Step 5: Node E added, 3 red links added. Node C now receives 3 harm edges (from B and E) and only 2 help edges (from B and D), so $\frac{3}{5}=0.6>\eta=0.5$. C is marked for removal (assuming its age exceeds $L_{S}$).
Step 6: Node C (and all its incident edges) removed. The network continues with nodes A, B, D, E.
2.1 Eigen Analysis of Adjacency Matrix
For additional insights into the structure of this adaptive network, we turned to Eigen Analysis which provides a higher level description of the simulation from the lower-level data. [25]
At each simulation tick, we generate an adjacency matrix by flattening the MultiGraph to a weighted graph by summing the number of edges: +1 for help edges and -1 for harm edges. Recall that at most 6 edges of one type (help or harm) can form between two nodes.
For example, in the above Figure 1 sample simulation at Step 5 (nodes A, B, C, D, E), the signed adjacency matrix $\mathbf{A}_{t}$ (where +1+1 denotes a help edge and 1-1 a harm edge, summed over parallel edges) is:
[…]
Entry $\mathbf{A}_{ij}$ indicates the net interaction from node $i$ to node $j$. For instance, $\mathbf{A}_{CA}=+1$ reflects C’s net helpful influence on A (2 help edges minus 1 harm edge).
We then approximate the dominant eigenpair (eigenvector and corresponding eigenvalue) $(\hat{\lambda}_{1},\mathbf{v}_{1})$ using a fixed-iteration power method with normalization at each step and a Rayleigh-quotient estimate,
$\hat{\lambda}_{1}\;=\;\frac{\mathbf{v}_{1}^{\top}\mathbf{A}_{t}\mathbf{v}_{1}}{\mathbf{v}_{1}^{\top}\mathbf{v}_{1}},$
terminating after at most 20 iterations or earlier if successive Rayleigh estimates satisfy $|\hat{\lambda}_{1}^{(k)}-\hat{\lambda}_{1}^{(k-1)}|<10^{-6}$.
To obtain a computationally efficient approximation of the sub-dominant eigenvalue, we perform a rank-one deflation by subtracting the estimated dominant mode,
$\mathbf{B}_{t}\;=\;\mathbf{A}_{t}\;-\;\hat{\lambda}_{1}\,\mathbf{v}_{1}\mathbf{v}_{1}^{\top},$
and reapply the same iterative procedure to $B_{t}$ to estimate $\hat{\lambda}_{2}$. We then report the spectral gap as $|\hat{\lambda}_{1}|-|\hat{\lambda}_{2}|$ and track the principal-angle change between successive dominant eigenvectors as a proxy for turnover in the network core. Because $A_{t}$ is generally non-symmetric, these spectral quantities should be interpreted as efficient heuristics for cohesion and regime change rather than as exact spectral characterizations at each time step.
The principal eigenvector for the adjacency matrix encodes the principal or dominant direction of flows or activity in the network. Its corresponding principal eigenvalue gives a measure to the ’strength’ of that dominance. An increasing steady principal eigenvalue also means that average connectivity (and so the potential for feedback/cooperation or harm) is still growing – the graph hasn’t saturated yet.
Our second measure, the spectral gap, is the difference between the two leading eigenvalues: $|\hat{\lambda}_{1}|-|\hat{\lambda}_{2}|$. A large and growing spectral gap implies there is one cohesive, resilient strongly connected component. In other words, the principal direction of connectivity has ’pulled away’ from any secondary directions. A small or falling gap would suggest there are many communities and more potential for fragmentation. Put another way, $\lambda_{1}$ tells us how dense the connections are, the spectral gap tells us how internally cohesive that density is.
Our third measure, the principal angle, reveals how stable the ’identity’ of the dominant nodes is. It measures the cosine similarity between two successive network states:
$\theta_{t}=\arccos\!\left(\frac{\mathbf{v}_{t-1}^{\top}\mathbf{v}_{t}}{\|\mathbf{v}_{t-1}\|\,\|\mathbf{v}_{t}\|}\right).$
A small, steady angle signals a persistent hub has formed; sharp increases imply regime shifts where new species take over the ‘core’ of the network.
Summary:
  • 1.
    High eigenvalue = high density and definition of help/harm interactions
  • 2.
    Principal angle = measures structural change between network states.
  • 3.
    Spectral gap = indicates dominance and stability of primary interaction pattern.
What follows are the results from several simulations. When parameters are altered, they are specifically listed below the figure.
3 Experiments & Results
In a systematic way, we explore how shifting the parameters (lifespan, harm-to-help ratio, binding chance) influence the emergence and persistence of strongly connected components (SCCs) in our directed network. Our primary interest was the evolution and robustness of cooperative structures (large SCCs) in networks characterized by varying degrees of antagonism (harm-to-help ratio). Eigen Analysis was performed to give a higher level characterization of the network’s behavior in the short- and long-run.
We enumerate and briefly explain the results found in the figures below:
  1. 1.
    Number of Nodes and Edges: At each time step $t$, the number of nodes and edges present in the graph. A node is added at every step with its own traits. Edges form based on the threshold rule. As it is the case with many network growth algorithms, existing nodes can create edges only with new nodes. Still, cycles can be formed with the proposed approach.
  2. 2.
    Fraction of Nodes in an SCC: The fraction of nodes at time $t$ in the graph that are part of a strongly connected component.
  3. 3.
    Fraction of Nodes in largest SCC: The fraction of nodes at time $t$ in the graph that are part of the largest strongly connected component.
  4. 4.
    In-Degree of Harm and Help Edges: The distribution of the number of inbound help and harm edges at the end of the simulation. For figures 1-3, after 8,000 steps.
  5. 5.
    Out-Degree of Harm and Help Edges: The distribution of the number of outbound help and harm edges at the end of the simulation. For figures 1-3, after 8,000 steps.
  6. 6.
    Mean Dominant Eigenvalue: Measure of the density of connections in the direction of the principal eigenvector.
  7. 7.
    Absolute Mean Dominant Eigenvalue: Removing the sign of the above reveals just the intensity of any shifts in the principal eigenvector, positive or negative.
  8. 8.
    Principal Angle: The angle between the principal eigenvectors between network states.
  9. 9.
    Spectral Gap: The difference between the first and second principal eigenvalues: $|\hat{\lambda}_{1}|-|\hat{\lambda}_{2}|$.
3.1 Harm-To-Help Ratio and Lifespan Dynamics
3.1.1 Low antagonism ($\rho$ below 0.5~0.5)
When cooperative interactions dominate (harm ratio $\rho<0.5$), we observed exponential growth in directed edge formation, resulting in a linearly increasing SCC size relative to the number of nodes in the network. This scenario characterizes a cooperative and expansive community where nodes not only persist, but integrate rapidly into a collective structure. Once such a community emerges, with a majority of help edges, it would be difficult for its elements to be removed.
Figures 3,4,5 show the results of the simulation from solely increasing the Harm-To-Help ratio from $\rho=0.3,0.5,0.7$ with all other parameters fixed. At first, we can restrict our view to the differences in the number of nodes and edges at step $t$, the largest SCC size and largest SCC fraction.
3.1.2 High antagonism ($\rho$ above 0.5~0.5)
For harm-to-help ratios exceeding 0.5, a strongly connected community still emerges and persists, provided nodes’ lifespan surpasses $\approx 10$ timesteps. Below this number, an SCC will fail to form reliably if $\rho$ is sufficiently high. See Figure 2.
Figure 2: Above harm ratio $\rho>0.5$, lifespan below $L=10$ will not reliably produce an SCC.
Notably here, SCC size does not grow indefinitely; rather, it reaches an upper bound and fluctuates around a steady-state maximum. This equilibrium state depends directly on both lifespan and the binding chance parameters. Increasing antagonism (harm ratio) gradually reduces the steady-state size of the SCC, yet even with extremely antagonistic environments (harm ratio approaching 0.9), some significant level of persistent inter-connectedness remains, as tenuous as it is.
Figure 3: Harm-To-Help: 0.3, Lifespan = 100, Bind-Chance = 0.003
Figure 4: Harm-To-Help: 0.5, Lifespan = 100, Bind-Chance = 0.003
Figure 5: Harm-To-Help: 0.7, Lifespan = 100, Bind-Chance = 0.003
3.2 Influence of Binding Chance
Adjusting the binding chance, $p_{max}$ which scales the likelihood threshold for interactions, primarily impacts the rate of convergence toward the final dynamic state (Figure 10). Increasing this parameter leads more quickly to equilibrium behaviors without substantially altering the qualitative behavior of SCC growth.
However, in scenarios where the harm ratio exceeded 0.5, a higher binding chance increased the fraction of nodes in the SCC, but overall slightly decreased the number of nodes in the equilibrium size of the SCC. This suggests that increasing the likelihood of interaction partially counteracts the destabilizing effect of widespread antagonism. See Figure 6.
$p_{max}$ Fraction in SCC Final SCC Size
0.005 $\approx$0.8 $\approx$210
0.0075 $\approx$0.9 $\approx$200
0.015 $\approx$0.98 $\approx$175
0.02 $\approx$1.0 $\approx$150
Figure 6: 30,000 Steps, $L_{S}=100$, $\rho=0.7$.
3.3 Node Age and Persistence
We noted a ‘demographic’ shift among highly-connected nodes (in the top 10% of in-degree) as antagonism increased. Specifically, and perhaps not surprisingly, networks with higher antagonism were increasingly dominated by younger nodes. High-harm environments shorten the node survival significantly, reducing the presence of “older” and more established nodes. Despite this shorter persistence, large SCCs remain viable and indicate a dynamically renewing structure of cooperation, which is resilient amidst high node turnover.
When viewing log(age) vs log(degree) we see bands or clustering when Harm-To-Help is <0.5<0.5 which dissipates and trends towards younger nodes as the network is stressed by harmful interactions. See Figure 7.
Figure 7: Log(age) vs Log(degree) after completing 10,000 steps of the simulation. Each point represents a node in the network at N=10,000 where $p_{max}=0.003$, at varying levels of harm: $\rho=0.2,0.5,0.8$. See video animation link for time-lapse visualization.
4 Discussion
Clarifying Scope of Vanilla Model
The framework introduced here operates deliberately at a higher level of abstraction. Rather than modeling specific biochemical, ecological, or evolutionary mechanisms that generate interactions (e.g., metabolic exchange, spatial structuring, gene loss, or adaptive specialization), the model treats the interaction network itself as the primary object of study. In this sense, it explores the space of possible help–harm configurations that such mechanisms may produce, rather than the mechanisms themselves. This abstraction allows us to identify network-level regularities—such as the persistence of large strongly connected components under mixed antagonism and support—that are potentially shared across microbial, economic, and other complex systems. At the same time, this perspective necessarily leaves unresolved questions about how particular interaction patterns arise, how they are modulated by spatial or metabolic constraints, or how they evolve over long timescales. The present results should therefore be interpreted as constraints on, rather than substitutes for, more detailed mechanistic models: any system whose interaction structure falls within this class would be expected to exhibit similar collective behavior, regardless of the underlying generative process.
4.1 Robustness of SCCs under High Antagonism
Our model demonstrates robustness of large SCC structure even when antagonistic interactions (harmful edges) significantly outnumber supportive interactions. This finding could be an extension of the concept of ‘islands of cooperation’ in [21]. Specifically, even with high harm-to-help ratios ($\rho>0.5$), SCCs consistently emerge and persist at steady-state equilibria. This implies the presence of inherent resilience within the network structure, which allows it to dynamically reconfigure to maintain cooperative viability despite continuous antagonistic pressure. The results on node “age” vs degree distribution lend another perspective on the long run dynamics, which we discuss later.
The interaction structure considered here assumes that benefits and harms are effectively public, an assumption that simplifies analysis but departs from many microbial systems in which goods are partially privatized or retained to varying degrees. In natural communities, metabolites may be incompletely shared, spatially constrained, or dynamically regulated, producing interaction strengths that vary over time and context.
The stability of these SCCs under high antagonism can be explained in network-theoretic terms: while harmful edges break node support structures, nodes with multiple positive inbound edges form resilient core communities, even when positive edges have a low probability of being formed. This leads to a dynamic steady state characterized by continuous recruitment and turnover of peripheral nodes, rather than collapse of the entire structure. These results align conceptually with robustness observed in dynamic real-world networks under external stresses, pointing to perhaps a broader principle of cooperative resilience inherent to certain threshold-based network structures.
4.2 Thresholds and Regime Shifts
A central outcome of our simulations is the emergence of a regime change near the harm-to-help ratio $\rho_{c}\approx 0.6$. (Figure 8, 9) For values of $\rho$ below this threshold, the strongly connected component (SCC) grows steadily with the total node population, reflecting a regime of sustained interaction and widespread mutual support. Above this threshold, however, SCC growth transitions from indefinite expansion to a bounded, fluctuating steady state.
This transition is evident in the late-time population growth rate, which decreases as $\rho$ increases and crosses zero near $\rho_{c}\approx 0.6$ (Figure 8). We define the late-time growth rate $g$ as the average per-step change in population size over the final $W=500$ time steps:
$g\;=\;\frac{|V_{T}|-|V_{T-W}|}{W}.$
A corresponding collapse in the mean network degree is shown in Figure 9, indicating a loss of interaction density that accompanies the change in growth dynamics.
Figure 8: The late-time (last 500 steps) population growth rate $g$ as a function of the harm-to-help ratio $\rho$. For small $\rho$, the system exhibits sustained positive growth. As $\rho$ increases, the growth rate decreases and crosses zero near $\rho_{c}\approx 0.6$, marking a shift between distinct dynamical regimes, from sustained growth to a bounded population regime.
Figure 9: Mean total degree (in + out) as a function of the harm-to-help ratio $\rho$. As $\rho$ increases, the average connectivity of the network decreases, with a marked reduction occurring near $\rho_{c}\approx 0.6$. This reduction in interaction density coincides with the crossover from sustained to bounded population growth observed in Figure 8, suggesting the loss of sufficient interaction capacity underlies the shift between dynamical regimes in population behavior.
While the harm-to-help ratio $\rho$ determines whether the system exhibits sustained or bounded growth, the binding chance $p_{max}$ governs the scale at which global coordination emerges. Figure 10 shows the fraction of nodes in the largest strongly connected component as a function of population size for different values of $p_{max}$. Increasing $p_{max}$ reduces the population scale required for SCC formation, indicating a change in coordination regime onset rather than final network structure.
This behavior is reminiscent of the connectivity thresholds in percolation and random graph models, in which a “giant” component becomes likely once the effective connection density is sufficiently high. [8] In our finite, stochastic setting the onset is rounded, but increasing $p_{max}$ consistently shifts coordination to smaller population sizes. Together, these results indicate that $\rho$ governs the growth–boundedness regime, whereas $p_{max}$ governs the population scale at which coordinated SCC structure typically emerges.
Figure 10: Onset of the largest strongly connected component (SCC) as a function of population size for varying bind chance $p_{max}$ ($\rho=0.2$, $L_{S}=100$). Increasing $p_{max}$ shifts the emergence of global connectivity to progressively smaller population sizes, indicating a transition in the onset of coordination. All curves converge to a near-unity SCC fraction, showing that $p_{max}$ primarily controls the timing of collective organization.
These thresholds offer concrete and testable predictions about the conditions required for cooperative network formation and may be connected to critical phenomena observed in physical and complex systems.
4.3 Eigen Analysis
Eigen Analysis of the network’s adjacency matrix provided us more insight into structural and dynamical features. The principal eigenvalue (approximated by method outlined in Section 2.1) measures overall connectivity strength along the dominant eigenvector direction, which reflects the magnitude and stability of cooperation (or antagonism) in the network. Higher eigenvalues indicate denser and strongly coupled cooperative or competitive interactions. In our simulations, rising mean eigenvalues (observed with $\rho<0.5$ ) corresponded to increasingly stable and interconnected SCCs.
The spectral gap (difference between the two largest eigenvalues) serves as a helpful secondary indicator, measuring how distinctly the principal ‘mode’ of interaction separates from other competing modes. A widening gap implies a clearly dominant, cohesive community structure, whereas narrowing gaps suggest fragmentation, multiple competing SCCs, or potential instability. Our results consistently showed that networks with high cooperation exhibit both large eigenvalues and pronounced spectral gaps, reaffirming the stability and dominance of a singular cooperative community structure.
Finally, the principal angle between consecutive eigenvectors quantifies the stability of core network structure across simulation steps. Small and stable principal angles indicate persistent dominance by specific nodes or clusters, while sudden spikes highlight structural rearrangements or shifts in network “leadership.” This measure gives additional, nuanced insights into temporal stability and dynamic structural shifts within the network. Consistent with our other measures, when the harm ratio was high, the principal angle fluctuated with increasing magnitude and signaled substantial shifts of influence of the key nodes in the support structure.
4.4 Age Distribution and Community Stability
The interplay between node lifespan and network antagonism produces distinctive patterns in node age (measured in time - birth-step) distributions (Figure 7). Under conditions of low antagonism (low $\rho$), older nodes with high connectivity become established, reflecting stable core communities. Conversely, increased antagonism rapidly shifts this balance toward younger nodes, as frequent node removal due to harmful interactions reduces node longevity. Interestingly, despite the reduction in older, established nodes, SCCs remain viable by continuously replenishing their core through newly added nodes. This activity reflects a structurally dynamic equilibrium rather than a fixed, static composition. Investigating the details behind the bands and clusters we see in the logarithmic plot may yield more insights. It is notable that at $\rho=0.5$ we see a cluster of nodes that are persisting most of the simulation, but have a reduced overall degree.
This age-based dynamic provides additional perspective on network stability and resilience. The continual integration of younger nodes suggests that organization under antagonism relies not on individual node longevity, but the network’s capacity for rapid turnover and ongoing renewal. This could mirror adaptive responses observed in other complex systems where dynamic reconfiguration rather than static preservation promotes robustness.
4.5 Binding Range as Ecological Analog
The binding range parameter, which can be thought of as the probability or “ease” of forming interactions (akin to biochemical or ecological compatibility), accelerates the dynamics of community structuring. It is interesting this parameter does not radically alter long-term outcomes but modulates the speed and slightly reduces the SCC’s ceiling under antagonistic conditions. In ecological contexts, this could suggest ecosystems with higher interaction potentials (dense signaling, physical proximity, etc) might achieve stable cooperation more rapidly and sustain slightly larger cohesive structures despite competitive pressures. Nevertheless, even in cases with a low binding range, it is “just a matter of time” before cooperative communities (autocatalytic sets, ecosystems (biological, economic, cultural), etc.) will be formed.
4.6 Emergent Levels of Organization in Diverse Systems
Our adaptive threshold network model, while originally inspired by microbial community dynamics, embodies general principles of emergent cooperation and the formation of higher-level organizational structures. The minimal assumptions we built into the model make it applicable to a wide variety of complex systems.
The dynamics observed here resonate strongly with other computational works  [1] and the concept of collective affordance sets [18]: where diverse system elements spontaneously combine their causal properties to create new functional wholes. Our findings illustrate clearly that, given sufficient time and diversity of interactions, these self-sustaining organizational structures are not merely possible but can be robustly expected to emerge.
In economic networks, analogous dynamics are not hard to find. Various goods, services, technologies, and businesses interact through interdependent relationships which support or undermine one another and give rise to intricate webs of economic cooperation. These adaptive networks continuously reorganize as new elements appear and obsolete ones exit, mirroring our model’s node addition and culling dynamics. Through this lens, our model can offer quantitative insights into the conditions under which economic ecosystems reliably generate new collective structures, whether clusters of industry specialization, innovation hubs, or resilient supply chains.
Similarly, ecosystems more broadly demonstrate these principles. Rich biodiversity provides a palette of interactions, allowing ecosystems to self-organize into resilient structures such as trophic networks, mutualistic partnerships, and adaptive community assemblies. Our work aligns with classical general systems theory [29] and recent empirical studies demonstrating self-organizing structures in ecological contexts, such as microbial communities [11] and biofilms [13].
Biofilms represent another powerful biological parallel. Individual bacterial species, each with unique traits and potentially antagonistic interactions, reliably assemble into complex multicellular aggregates. The emergent properties of biofilms such as shared resource utilization, collective defense mechanisms, and coordinated behavior, demonstrate exactly the kind of spontaneous higher-level organization our model predicts. Indeed, such structures can be viewed as an evolutionary step toward multicellularity [13, 7, 5].
Placing our model within the broader framework of major evolutionary transitions [20, 19] also yields valuable insights. It highlights conditions under which Darwinian preadaptations can coalesce into higher-order systems, suppressing or transcending individual-level competition. So, our results offer a theoretical foundation for understanding transitions from simple chemical reaction networks to living cells, from unicellular organisms to multicellular structures, from individuals to groups [14], and from isolated economic agents to integrated economies.
By demonstrating robust self-organization through minimal assumptions, our adaptive threshold network model not only sheds light on microbial dynamics but provides a versatile conceptual tool applicable across biological, economic, and social domains. Given enough time, diversity, and interaction potential, the emergence of cooperative structures and new levels of organization appears to be not merely possible but inevitable.
5 Conclusions and Future Work
We have demonstrated that, under minimal assumptions, adaptive threshold networks naturally give rise to communities of support and cooperation, even in the presence of substantial antagonism. The primary insight is that the vast diversity of possible interactions enables complex systems to reliably find paths towards sustained cooperation. The two parameter-dependent regime shifts identified in the model suggest that at some point quantity turns into quality. Thus, our results provide support for the conjecture that life inherently ”finds a way,” tinkering to build cooperative structures despite adverse conditions.
Our current model intentionally excludes stronger forms of evolutionary and co-evolutionary dynamics, setting these aspects as key areas for future development. Integrating evolutionary games — which have been widely explored elsewhere [20, 22, 26] — could enrich our understanding of how cooperation not only emerges but is subsequently refined and stabilized through selective pressures.
Another aspect of our model that remains to be explored is the fact that it can be seen as an example of strong emergence [2, 3, 27, 16] and downward causation [6, 4, 10, 12]. Once communities are formed (at a higher scale), they influence the survivability of nodes (at a lower scale). Thus, it is not possible to reduce the behavior of the system to the properties of its elements. Nevertheless, there are epistemological nuances that should be considered and are beyond the scope of this paper.
Furthermore, the notion of resources is presently implicit in the parameter ‘lifespan,’ which dictates how long a node persists without adequate support. Introducing explicit resource dynamics and spatial constraints would enhance biological realism, enabling the exploration of more nuanced ecological and economic scenarios. Spatial structure, in particular, can affect opportunities for interaction and substantially impact emergent dynamics. [17, 23]
A natural step is to identify biologically and economically realistic parameter ranges, particularly for lifespan and interaction (binding) probabilities. Lifespan can affect the short- and long-run behaviors of the network, preventing formation if set too low, or inflating the size of the SCC as it increases. We expect these steady-states can be analytically approximated by functions of lifespan and binding probabilities, which is an open avenue for future work. Another investigation would be to what degree lifespan is emphasized here as one of our minimal assumptions. For example, a species’ death may continue to provide an affordance (easier access to food, etc.) but these interactions would be lost in our model.
We can assume that evolution will favor a balanced $\rho$ ratio: too much harm is not sustainable. (This perspective is consistent with empirical studies of Streptomyces communities in soil, where dense networks of antagonistic and growth-promoting interactions coexist, exhibiting balanced interaction probabilities, strong reciprocity, and rapid evolutionary turnover rather than static equilibrium [28].) However, only help might hinder adaptability and evolvability  [15]. To explore this hypothesis systematically, we plan to introduce evolutionary feedback explicitly in the next iteration of our model. Edge weights will dynamically evolve, strengthening or weakening interactions probabilistically over time, reflecting real-world processes of mutual adaptation and co-evolution. Preliminary results and theory suggest this enhancement should lead to rich dynamics, potentially revealing other novel transitions or critical thresholds that influence community resilience and connection.
Finally, empirical validation remains paramount. We eagerly await the experimental outcomes from Jan Dijksterhuis’ ongoing microbial experiments involving 140 microbial species. These empirical findings may yet validate our model’s assumptions but also inspire refinements. Will we find evidence of persisting networks of support? Will these networks look the same in different soil plots? Whatever daylight exists between our findings and the empirical results will motivate, or afford, more interesting explorations.
Acknowledgements
We thank Andrea Roli and Jérome Michaud for their insightful comments and valuable feedback, which greatly improved the manuscript.